# Copyright Infringement Detection in Text-to-Image Diffusion Models via Differential Privacy

### v2

@@ 1 Introduction
Recent advances in large vision models have improved the realism of image synthesis. While they are widely adopted in creative industries and public platforms, they have also raised serious concerns over copyright infringement. Researches (Somepalli et al., 2023; Cilloni, Fleming, and Walter, 2023; Carlini et al., 2023) find that models such as Stable Diffusion (Rombach et al., 2022) may memorize and reproduce contents in training datasets, including those with unclear permission or copyright violation. These risks are exacerbated by the lack of transparency regarding the provenance of training data in most models. As a result, there is an urgent need for accurate and reliable post-hoc methods to detect potential copyright infringements in models.
Recent studies have attempted to solve this issue. For example, CopyScope (Zhou et al., 2023) proposes a model-level framework to quantify each component’s contribution to potential copyright infringement in diffusion workflows, by evaluating FID-based similarity and Shapley value of attributing responsibility. However, it does not identify specific infringed concepts or samples, and therefore cannot provide concrete legal evidence of infringement. Other works (Wang et al., 2024; Xu et al., 2025) detect infringement via prompt queries or prompt engineering work. However, prompt-based detection is inherently fragile, as it depends on constructing specific conditioning prompts to trigger infringing outputs, which can be easily affected by minor changes, including model updates and sampling randomness. This limits generalization across models and datasets. It also lacks interpretability, as the underlying causes of infringement remain unclear, and reproducibility, due to stochastic generation and non-deterministic model behavior.
@@ 2 Background and Related Work
To address these challenges, several regulatory frameworks (NIST, 2023; State Internet Information Office of the People’s Republic of China et al., 2023; Official Journal of the European Union, 2024; Oliver et al., 2025) are being introduced globally. However, they are difficult to enforce in practice due to the opacity of large-scale models. Once a model has been trained, it is impractical to retrain from scratch to satisfy these regulations. Hence, it becomes vital to determine whether a specific data point has been memorized and is urgently needed for post-hoc methods to detect and remove unauthorized content from deployed models.
On the technical front, Zhou et al. propose CopyScope—a model-level framework that quantifies copyright infringement in the full diffusion workflow in three stages: identify influential components, quantify by FID-Shapley and evaluate contributions of models. DIAGNOSIS (Wang et al., 2023) detects the marked copyright infringement by first coating the protected dataset, then approximating the memorization strength, and making a hypothesis testing. Ma et al. provides a dataset and benchmark for copyright infringement protection in data unlearning of text-to-image diffusion models. It introduces a metric that evaluates similarity between two images from both semantic and stylistic perspectives.
In the copyright infringement detection of LVLMs, Wang et al. propose a prompt engineering method that generates prompts to trigger IP violations in Large Language Models (LLMs) under black-box settings. Xu et al. introduce an IP benchmark dataset and find out that LVLMs tend to misclassify the negative IP samples using in-context learning. Chiba-Okabe and Su design a systematic evaluation criterion to quantify and estimate the originality level of data, and introduce PREGen to modify outputs and lower originality values.
@@ 3.1 Diffusion Models
Diffusion Models
Diffusion Models (DMs) (Ho, Jain, and Abbeel, 2020) are latent variable generative models which learns a data distribution $p(x)$ by gradually denoising variables sampled from a Gaussian distribution. This process is achieved by reversing a fixed Markov chain of stochastic noise injection. Formally, the forward process gradually adds noise to a data sample $x_{0}\sim q(x_{0})$ over $T$ steps through a sequence of Gaussian transitions:
@@
Latent Diffusion Models
To reduce model complexity and preserve the image details, Latent Diffusion Models (LDMs) (Rombach et al., 2022) employ the DM training in the lower-dimensional latent space $z$ of an encoder $\mathcal{E}$, and the images $\tilde{x}$ can be reconstructed by the decoder $\mathcal{D}(z)$, giving $\tilde{x}=\mathcal{D}(z)=\mathcal{D}(\mathcal{E}(x))$. The training objective of an LDM can be written as:
$\mathcal{L}_{\text{{LDM}}}=\mathbb{E}_{\mathcal{E}(x),\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t)\right\|_{2}^{2}\right].$ (3)
@@ 3.2 Unlearning of Diffusion Models
Machine unlearning (Bourtoule et al., 2021) is a task of removing the influence of specific dataset $A$ from a model trained on $X$ without retraining from scratch. In DMs, Naive Deletion fine-tunes the model on the retained dataset $X\setminus A$ by minimizing the simplified evidence-based lower bound (ELBO); NegGrad (Golatkar, Achille, and Soatto, 2020) performs gradient ascent on the unlearning set $A$, while Subtracted Importance Sampled Scores (SISS) (Alberti et al., 2025) is a novel objective that interpolates between Naive Deletion and NegGrad using importance sampling, allowing efficient and stable fine-tuning with theoretical guarantees.
@@ 3.3 Differential Privacy
Differential privacy (Dwork et al., 2006) is a formal notion of algorithmic privacy, which aims to prevent the release of private information.
@@ 4 Copyright Infringement
Copyright infringement occurs when an algorithm $G$ (e.g., a text-to-image model) is trained on a dataset $D$ that includes a subset of copyrighted or unauthorized data samples $D_{C}=\{x_{c_{1}},x_{c_{2}},\dots,x_{c_{n}}\}\subset D$, without permission. Evidence for such an infringement claim is typically established by demonstrating that the model’s outputs reproduce or are substantially similar to the protectable expressive elements of the samples in $D_{c}$.
To establish a theoretical basis for analyzing copyright infringement, we formalize differential privacy (Bousquet, Livni, and Moran, 2020) in this context:
@@
This approach provides a principled means to quantify the risk of copyright infringement.
4.1 Formalization of Differential Privacy
Conditional Publicity for Generative Models
Previous researches have revealed significant privacy vulnerabilities in the outputs of diffusion models. For instance, membership inference (Cilloni, Fleming, and Walter, 2023) has achieved success rates exceeding 60%; data extraction and reproduction (Somepalli et al., 2023; Carlini et al., 2023) have shown that diffusion models can memorize and reproduce individual training images. Building on these findings, we hypothesize that diffusion models exhibit almost no conditional differential privacy, but with much more publicity:
Definition 3 (Conditional Publicity for Diffusion Models).
$\gg\Pr[G(\theta_{D^{\prime}},p)\in S]$
where $\theta_{D}$ and $\theta_{D^{\prime}}$ are the model parameters obtained by training model $G$ on datasets $D$ and $D^{\prime}$ respectively. Here, $\epsilon>200$ indicates a substantial violation of privacy, i.e., the model output strongly depends on the presence of a specific concept in the training set.
Relative Privacy for Non-Infringed Data
The notion of conditional differential publicity implies that a generative model could significantly alter its relevant outputs when a single concept in the training dataset is modified. On the contrary, an input that the model has not seen, namely, non-infringed data, could be relatively private:
@@ 4.1 Formalization of Differential Privacy
Let $x$ be a data point (e.g., a prompt-image pair) that does not appear in any subset of the training dataset $D$ or its neighbouring dataset $D^{\prime}$. Then, under Def. 2, the generative model $G$ satisfies $(0,0)$-conditional differential privacy with respect to $x$ for any prompt $p$.
@@
The above proposition and definition formalize the notion that non-infringed data should not exert measurable influence on the model’s output distribution. In contrast to infringed data that violates conditional privacy guarantees, non-infringed data satisfies strict or approximate relative privacy. This theoretical distinction lays the foundation for our detection criterion: if a concept induces statistically significant behavioral change in model outputs, it suggests potential memorization and copyright infringement.
4.2 Definition of Copyright Infringement and Non-Infringement
Building upon the formalism of conditional differential privacy and publicity, we now define the mathematical criteria for: when a generative model $G$ can be considered to have infringed upon a specific data point or concept (e.g., a copyrighted image or IP), and when it does not, based on the model behavior.
Definition 5 (Copyright Infringement).
$\Pr[G(\theta_{D},p)\in S]\gg\Pr[G(\theta_{D^{\prime}},p)\in S],$ (12)
where $D^{\prime}=D\setminus\{x_{c}\}$ is a neighboring dataset.
Definition 6 (Copyright Non-Infringement).
$\Pr[G(\theta_{D},p)\in S]=\Pr[G(\theta_{D^{\prime}},p)\in S].$ (13)
To allow for a relaxed setting, we say that $G$ satisfies approximate non-infringement if the following $(\epsilon,\delta)$-differential privacy holds:
@@ 4.2 Definition of Copyright Infringement and Non-Infringement
This definition ensures that the output distribution is invariant with respect to the inclusion or exclusion of $x$, thereby approximately satisfying $(0,0)$-conditional differential privacy and guaranteeing that $x$ does not influence model behavior. Taken together, these definitions provide a theoretical framework for copyright infringement detection.
5 Detection of Copyright Infringement
5.1 Definition of Copyright Infringement Detection
Definition 7 (Detection of Copyright Infringement).
5.2 Conditional Sensitivity Metric
5.3 D-Plus-Minus (DPM) Detection
Problem Setting
Detection of copyright infringement faces several practical challenges, such as scalability, inaccessibility of training data, conditional input unavailability, and insufficient theoretical guarantees (detailed analysis of challenges available in Appendix A). In light of these issues, our work operates under a realistic and challenging set of assumptions: (1) white-box access to a pretrained model; (2) absence of corresponding input prompt; (3) inaccessibility of training data. Given these constraints, we aim to identify evidence of infringement as reflected in the model’s observable behavior, grounded in the above theoretical framework. Figure 2 illustrates the core procedure of our proposed D-Plus-Minus detection framework. It contains several steps: concept extraction, image collection, branch training, conditional sensitivity measurement, statistics analysis, and branch merging.
Pre-Processing: Concept extraction and Image Collection
Branch Training
Conditional Sensitivity Measurement
Here, $M(\cdot)=\text{CLIP}_{\mathcal{I}}(\cdot)$ denotes the CLIP (Radford et al., 2021) image encoder, which serves as a query function to capture the semantics similarity between two outputs. The nearer the conditional sensitivity is to $1$, the less sensitive it is.
Statistics Analysis
Branch Merging
$\quad\text{DPM}(M,\hat{x}_{i},D_{\text{total}})$
6 Hierarchical Categories of Copyright Infringement and CIDD Dataset
To comprehensively categorize copyright infringement in generative models, we propose a hierarchical taxonomy containing four levels of content resemblance, as shown in table 1. While low-level technical features, such as edge detectors, texture filters or color histograms, may cause perceptual similarity, they are typically insufficient on their own to constitute legal infringement. Instead, it is more likely to be substantiated when higher-level similarities, such as character compositions or stylistic elements, indicate derivation from copyrighted works. The top-level semantics similarity is often shown in videos with a time sequence, rather than just one image. Details are listed in Appendix C.
    Level   Categories   Examples
  Level 1   Technics   —
  Level 2   Content   Human Face*
  Level 3-1   Structure   Architecture*
  Level 3-2   Style   Arts Painting*
  Level 4   Semantics   Plots & Themes
Table 1: Hierarchical Categories of Copyright Infringement. It is ordered from low-level perceptual features to high-level conceptual constructs. “*” means the used classes in CIDD dataset.

### v3

@@ 1 Introduction
Recent advances in large vision models have improved the realism of image synthesis. While they are widely adopted in creative industries and public platforms, they have also raised serious concerns over copyright infringement. Researchers (somepalli2023diffusion; cilloni2023privacythreatsstablediffusion; carlini2023extracting) find that models such as Stable Diffusion (rombach2022high) may memorize and reproduce contents in training datasets, including those with unclear permission or copyright violation. These risks are exacerbated by the lack of transparency regarding the provenance of training data in most models. As a result, there is an urgent need for accurate and reliable post-hoc methods to detect potential copyright infringements in models.
Recent studies have attempted to solve this issue. For example, CopyScope (zhou2023copyscopemodellevelcopyrightinfringement) proposes a model-level framework to quantify each component’s contribution to potential copyright infringement in diffusion workflows, by evaluating FID-based similarity and Shapley value of attributing responsibility. However, it does not identify specific infringed concepts or samples, and therefore cannot provide concrete legal evidence of infringement. Other works (wang2024evaluatingmitigatingipinfringement; xu2025largevisionlanguagemodelsdetect) detect infringement via prompt queries or prompt engineering work. However, prompt-based detection is inherently fragile, as it depends on constructing specific conditioning prompts to trigger infringing outputs, which can be easily affected by minor changes, including model updates and sampling randomness. This limits generalization across models and datasets. It also lacks interpretability, as the underlying causes of infringement remain unclear, and reproducibility, due to stochastic generation and non-deterministic model behavior.
@@ 2 Background and Related Work
To address these challenges, several regulatory frameworks (airmf2023; interimgai2023; euaiact2024; generalpurposeai2025) are being introduced globally. However, they are difficult to enforce in practice due to the opacity of large-scale models. Once a model has been trained, it is impractical to retrain from scratch to satisfy these regulations. Hence, it becomes vital to determine whether a specific data point has been memorized and is urgently needed for post-hoc methods to detect and remove unauthorized content from deployed models.
On the technical front, zhou2023copyscopemodellevelcopyrightinfringement propose CopyScope—a model-level framework that quantifies copyright infringement in the full diffusion workflow in three stages: identify influential components, quantify by FID-Shapley and evaluate contributions of models. DIAGNOSIS (wang2023diagnosis) detects the marked copyright infringement by first coating the protected dataset, then approximating the memorization strength, and making a hypothesis testing. ma2024datasetbenchmarkcopyrightinfringement provides a dataset and benchmark for copyright infringement protection in data unlearning of text-to-image diffusion models. It introduces a metric that evaluates similarity between two images from both semantic and stylistic perspectives.
In the copyright infringement detection of LVLMs, wang2024evaluatingmitigatingipinfringement propose a prompt engineering method that generates prompts to trigger IP violations in Large Language Models (LLMs) under black-box settings. xu2025largevisionlanguagemodelsdetect introduce an IP benchmark dataset and find out that LVLMs tend to misclassify the negative IP samples using in-context learning. Chiba_Okabe_2025 design a systematic evaluation criterion to quantify and estimate the originality level of data, and introduce PREGen to modify outputs and lower originality values.
@@ 3.1 Diffusion Models
Diffusion Models (DMs) (ho2020denoising) are latent variable generative models which learns a data distribution $p(x)$ by gradually denoising variables sampled from a Gaussian distribution. This process is achieved by reversing a fixed Markov chain of stochastic noise injection. Formally, the forward process gradually adds noise to a data sample $x_{0}\sim q(x_{0})$ over $T$ steps through a sequence of Gaussian transitions:
@@
To reduce model complexity and preserve image details, Latent Diffusion Models (LDMs) (rombach2022high) employ the DM training in the lower-dimensional latent space, mapped by an encoder and reconstructed by a decoder.
@@ 3.2 Unlearning of Diffusion Models
Machine unlearning (bourtoule2021machine) is a task of removing the influence of specific dataset $A$ from a model trained on $X$ without retraining from scratch. In DMs, Naive Deletion fine-tunes the model on the retained dataset $X\setminus A$ by minimizing the simplified evidence-based lower bound (ELBO); other approaches include NegGrad (golatkar2020eternal) and SISS (alberti2025dataunlearningdiffusionmodels).
@@ 3.3 Differential Privacy
Differential privacy (DP) (dwork2006calibrating) is a formal notion of algorithmic privacy, which aims to prevent the release of private information.
@@ 4 Copyright Infringement
Copyright infringement occurs when an algorithm $G$ (e.g., a text-to-image model) is trained on a dataset $D$ that includes a subset of copyrighted or unauthorized data samples $D_{C}=\{x_{c_{i}}\}\subset D$, without permission. A claim of infringement is typically established by showing that model’s outputs reproduce or are substantially similar to the protectable expressive elements of the copyrighted samples in $D_{c}$.
4.1 Formalization of Differential Privacy
Some researchers (cilloni2023privacythreatsstablediffusion; somepalli2023diffusion; carlini2023extracting) have revealed significant privacy vulnerabilities in the outputs of diffusion models. For instance, membership inference (cilloni2023privacythreatsstablediffusion) has achieved success rates exceeding 60%; data extraction and reproduction (somepalli2023diffusion; carlini2023extracting) have shown that diffusion models can memorize and reproduce individual training images. It suggests that these models exhibit almost no conditional differential privacy regarding the training dataset, but with more publicity:
Definition 2 (Conditional Publicity for Diffusion Models).
$\gg\Pr[G(\theta_{D^{\prime}},p)\in S],$
where $\theta_{D}$ and $\theta_{D^{\prime}}$ are the model parameters obtained by training model $G$ on datasets $D$ and $D^{\prime}$ respectively.
Here, $\epsilon>200$ indicates a substantial violation of privacy, i.e., the model output strongly depends on the presence of a specific concept in the training set.
From the perspective of differential privacy, a generative model’s relevant outputs could be drastically altered when a single copyrighted training concept is modified. Conversely, an input that the model has not seen, namely, non-infringed data, could be relatively private.
A detailed mathematical description of differential privacy in diffusion models is given in the appendix.
4.2 Definition of Copyright Infringement
Building upon the formalization of differential privacy, we now define the mathematical criteria for: when a generative model $G$ can be considered to have infringed upon a specific data point or concept (e.g., a copyrighted image or IP), and when it does not, based on the model behavior.
Definition 3 (Copyright Infringement).
$\Pr[G(\theta_{D},p)\in S]\gg\Pr[G(\theta_{D^{\prime}},p)\in S],$ (7)
where $D^{\prime}=D\setminus\{x_{c}\}$ is a neighboring dataset.
Definition 4 (Copyright Non-Infringement).
$\Pr[G(\theta_{D},p)\in S]=\Pr[G(\theta_{D^{\prime}},p)\in S].$ (8)
This definition ensures that the output distribution is invariant with respect to the inclusion or exclusion of $x$, thereby guaranteeing that $x$ does not influence model behavior. Taken together, these definitions provide a theoretical framework for copyright infringement detection.
5 Detection of Copyright Infringement
5.1 Definition of Infringement Detection
Definition 5 (Detection of Copyright Infringement).
5.2 Conditional Sensitivity Metric
5.3 D-Plus-Minus (DPM) Detection
Problem Setting
Detection of copyright infringement faces several practical challenges, such as scalability, inaccessibility of training data, conditional input unavailability, and insufficient theoretical guarantees. In light of these issues, our work operates under a realistic and challenging set of assumptions:
(1) white-box access to a pretrained model;
(2) absence of corresponding input prompt;
(3) inaccessibility of training data.
Given these constraints, we aim to identify evidence of infringement as reflected in the model’s observable behavior, grounded in the above theoretical framework. Figure 2 illustrates the core procedure of our proposed D-Plus-Minus detection framework. It contains several steps: concept extraction, image collection, branch training, conditional sensitivity measurement, statistics analysis, and branch merging.
Pre-Processing: Concept extraction and Image Collection
Branch Training
Conditional Sensitivity Measurement
Here, $M(\cdot)=\text{CLIP}_{\mathcal{I}}(\cdot)$ denotes the CLIP (radford2021learning) image encoder, which serves as a query function to capture the semantics similarity between two outputs. The nearer the conditional sensitivity is to $1$, the less sensitive it is.
Statistics Analysis
  Class SD1.4 SDXL-1.0 SANA-0.6B FLUX.1
Human Face 0.9011 0.8058 0.7011 0.6289 0.8062 0.7285 0.7531 0.6419
Merged Total 0.8071 0.6726 0.7800 0.7234 0.7914 0.6855 0.8257 0.7039
  Level Categories Examples
Level 1 Technics
Level 2 Content Human Face*
Level 3-1 Structure Architecture*
Level 3-2 Style Arts Painting*
Level 4 Semantics Plots & Themes
Table 2: Hierarchical Categories of Copyright Infringement. It is ordered from low-level perceptual features to high-level conceptual constructs. “*” means the classes in CIDD.
Figure 3: Qualitative visualization of two branches across different timesteps. Models tend to learn and unlearn faster with infringed samples, while slower on non-infringed ones, and cannot learn exact elements in the target images.
Branch Merging
$\text{DPM}(M,\hat{x}_{i},D_{\text{total}})$
6 Hierarchical Categories of Infringement
To comprehensively categorize copyright infringement in generative models, we propose a hierarchical taxonomy containing four levels of content resemblance, as shown in table 2. While low-level technical features, such as edge detectors, texture filters or color histograms, may cause perceptual similarity, they are typically insufficient on their own to constitute legal infringement. Instead, it is more likely to be substantiated when higher-level similarities, such as character compositions or stylistic elements, indicate derivation from copyrighted works. Furthermore, the highest level of semantics similarity is often shown in time-sequenced media (e.g., videos), rather than being confined to one image.
7 Copyright Infringement Detection Dataset
To overcome the limitations of existing datasets, we construct the Copyright Infringement Detection Dataset (CIDD). It contains several classes of orthogonal prompts and three image classes that are most likely to be infringed: human face, architecture, and arts painting, with a total of 429 concepts and 2,397 images. Each class is mapped to one hierarchical category in table 2. Crucially, CIDD includes both infringed and non-infringed concepts, each of which is annotated with a binary infringement label based on its source and content provenance, and is paired with 3 to 6 neighbourhood images, enabling robust learning and evaluation under weak and probabilistic assumptions.
Data examples, summaries, and collection methods are detailed in the appendix.
Figure 4: ROC curves in four representative models. The proposed DPM confidence score consistently outperforms individual branches in terms of AUC, demonstrating its superior capability in distinguishing infringed from non-infringed samples.
8 Experiment
As mentioned in Section 2, Copyscope cannot distinguish image-level infringement; DIAGNOSIS is an infringement detection method based on a priori dataset coating; CPDM can only evaluate the extent of infringement after data unlearning. These methods target unrealistic settings and different goals, making direct comparison with DPM inappropriate. As for detection in LLMs and LVLMs, DPM can also be extended to this setting, which we leave as a promising work in the future. In this section, we evaluate the proposed DPM on the CIDD dataset across four text-to-image models: Stable Diffusion v1.4 (SD1.4) (rombach2022high), Stable Diffusion XL (SDXL) (podell2023sdxlimprovinglatentdiffusion), SANA (xie2024sana), and FLUX (labs2025flux1kontextflowmatching) models.
We report both quantitative and qualitative results, and perform an ablation study to validate all modules in DPM. For the quantitative evaluation, we employ two standard metrics: ROC-AUC, which measures the ability to distinguish infringed from non-infringed samples, and Soft Accuracy (SoftAcc), a stricter metric that evaluates the numerical alignment between confidence scores and ground truth labels, penalizing slight deviations from the correct value.
8.1 Quantitative Experiments
We report per-class and aggregated results in Table 1. DPM consistently achieves a weighted-average AUC above 80% (as shown in Fig. 4) and SoftAcc above 72% across models, demonstrating its strong generalization.
Moreover, detection performance varies across classes, mainly driven by the distinct generalization capabilities of the CLIP models, the specific parameter configurations of the diffusion models, and the utilized fine-tuning timesteps.
8.2 Qualitative Experiments
To better understand the behaviors of our two-branch framework, we visualize representative samples in Fig. 3.
These results suggest that the infringed samples have triggered the model’s internal memory, causing it to reproduce memorized content with high fidelity; while the unlearning branch rapidly suppresses such resemblance, producing visually dissimilar outputs after just a few negative steps. For non-infringed samples, both branches tend to maintain stable generations with limited directional change.
8.3 Ablation Study and Robustness
We provide a detailed ablation study and robustness analysis in the appendix. We find that: (1) the proposed conditional sensitivity metric effectively distinguishes between infringed and non-infringed samples, as partially illustrated in Fig. 5; (2) merging of two branches and multiple timesteps both improve and stabilize detection performance; (3) image degradation affects little to the detection performance.
9 Conclusion
Appendix A Supplementary on the Theory of Copyright Infringement
A.1 Introduction of Differential Privacy
The notion of differential privacy: A randomized algorithm that receives a sequence of data points $\bar{x}$ as input is differentially private if removing/replacing a single data point in its input, does not affect its output $y$ by much; more accurately, for any event $E$ over the output $y$ that has non-negligible probability on input $\bar{x}$, then the probability remains non-negligible even after modifying one data point in $\bar{x}$.” — Synthetic Data Generators – Sequential and Private (bousquet2020synthetic)
Inspired by the quote above, the original Def. 1 can be reformulated for generative models as follows:
@@ A.1 Introduction of Differential Privacy
A.2 Relative Privacy for Non-Infringed Data
Based on the principle illustrated in Section 4.1, we formalize the following proposition:
@@ A.2 Relative Privacy for Non-Infringed Data
Let $x$ be a data point (e.g., a prompt-image pair) that does not appear in any subset of the training dataset $D$ or its neighbouring dataset $D^{\prime}$. Then, under Def. 6, the generative model $G$ satisfies $(0,0)$-conditional differential privacy with respect to $x$ for any prompt $p$.
@@
Similarly, while achieving strict non-infringement in Def. 4, as defined by $\Pr[G(\theta_{D},p)\in S]=\Pr[G(\theta_{D^{\prime}},p)\in S]$, is challenging in practice, we introduce a relaxed setting to account for minor distributional changes. This leads to the definition of approximate non-infringement, which leverages the concept of differential privacy:
Definition 8 (Approximate Copyright Non-Infringement).
A generative model $G$ satisfies approximate non-infringement if the following $(\epsilon,\delta)$-differential privacy holds:
@@
This definition ensures that the output distribution is invariant with respect to the inclusion or exclusion of $x$, thereby approximately satisfying $(0,0)$-conditional differential privacy and guaranteeing that $x$ does not influence model behavior.
Appendix B Copyright Infringement Detection Dataset (CIDD) Preview
Motivation